# CushionCatch: Compliant Catching Mechanism for Mobile Manipulators via Combined Optimization and Learning

### v1

@@ Abstract
This paper presents a framework to achieve compliant catching with cushioning mechanism(CCCM) for mobile manipulators. First, we introduce a two-level motion optimization scheme, comprising a high-level capture planner and a low-level joint planner. The low-level joint planner consists of two distinct components: Pre-Catching (PRC) planner and Post-Catching (POC) planner. Next, we propose a network that leverages the strengths of LSTM for temporal dependencies and positional encoding for spatial context(P-LSTM). P-LSTM is designed to effectively learn compliant control strategies from human demonstrations. To account for structural differences between humans and robots, safety constraints are incorporated into POC planner to avoid potential collisions. We validate the CCCM framework through both simulated and real-world ball-catching scenarios, achieving a success rate of 98.70% in simulation, 92.59% in real-world tests, and a 33.2% reduction in impact torques.
@@ I INTRODUCTION
Manipulating dynamic objects is challenging, requiring a sophisticated interplay of object detection, motion planning, and control [1]. In dynamic manipulation scenarios, tasks such as throwing [2, 3], juggling [4], and catching [5, 6] are essential. Catching dynamic objects requires quick and precise motion planning, followed by fast and accurate execution. For the motion planning of catching, the process can be divided into two phases, Pre-Catching (PRC) and Post-Catching (POC) [7]. In the POC phase, for a high-stiffness robot, the object may bounce away or generate significant impact forces upon contact. To address this, incorporating a cushioning mechanism during catching is beneficial, as it reduces impact forces [8, 9], prevents damage to both the robot and the object [10], and minimizes the risk of rebound, especially for high-speed or high-mass objects [11, 12]. Moreover, since humans naturally exhibit cushioning behavior when catching objects, an effective approach would be for the robot to learn this motion from humans.
Figure 1: The mobile manipulator’s motions under CCCM framework when catching a ball.
Mobile manipulators are particularly adept at dynamic manipulation tasks, outperforming conventional stationary robots [13, 14] with greater flexibility and efficiency [15, 16, 17]. Their mobility allows them to actively navigate and interact with objects [18, 19]. In catching tasks, most prior research focused on rigid control approaches, often overlooking the importance of compliant control [1, 6, 14, 20]. Mobile manipulators, with their wireless capabilities, are ideal for learning human-like compliant behavior. We focus on introducing a framework for compliant catching using human demonstration data, which is designed to be modular, allowing easy integration with other robotic systems.
Compared with previous works, the contributions of this letter are:
1. Propose a two-level optimization scheme that includes a high-level capture planner and a low-level joint planner. This scheme addresses all planning and tracking challenges during the POC and PRC phases, while also solving potential safety concerns arising from structural differences in imitation learning.
2. Propose a P-LSTM network, a sequence-to-sequence model that combines a Long-Short-Term Memory network(LSTM) with positional encoding to predict commands, enabling the effective learning of compliant control strategies from human demonstrations.
3. Simulation and physical experiments were conducted on a mobile manipulator to validate the effectiveness of the CCCM framework and the role of compliance. Catching success rates of 98.70% in simulation, 92.59% in physical experiments and a 33.2% reduction in impact torque were achieved.
@@ II Related Work
Many studies have examined ball catching with fixed-base manipulators [1, 6, 21, 22], focusing particularly on motion planning for the manipulator. Early research utilized heuristic methods to determine optimal grasping points and generated motion trajectories through interpolation. However, these methods often fall short of achieving the most effective capture motion of the robot. A more effective approach involves using nonlinear optimization to identify the optimal joint configuration for catching the ball, along with parameterization techniques such as five-order polynomials [20].
Imitation learning allows algorithms to learn from human demonstrations, simplifying the acquisition of complex behaviors [23]. [7] creates nucleated motion primitives from human demonstration data to generate compliant trajectories. [24] uses a Gaussian Mixture Model and a Long-Short Term Memory (LSTM) model to predict the trajectory of a manipulation. However, their approach does not account for the potential safety issues arising from structural differences between humans and robots.
Figure 2: Overview of the system framework: External sensors provide status for both the ball and the robot. The ball’s state is refined using a Kalman filter, which is then used to predict its future trajectory over a time horizon. The high-level capture planner computes the optimal capture state, $\boldsymbol{q}_{ca}$. The PRC planner generates rapid motion trajectories for $\boldsymbol{q}_{ca}$ and sends them to the joint position controller. The POC planner tracks the compliant velocity sequence from the P-LSTM network, which is designed to learn from human demonstrations.
@@ III-A System Model
The mobile manipulator we use includes a mobile base and a manipulator, with $n=n_{b}+n_{m}$ degrees of freedom. The state of the manipulator is described by joint positions $\boldsymbol{q}_{{m}}\in\mathbb{R}^{n_{m}}$. For a mobile base with a differential drive, we use the full location description $(x_{b},y_{b},\varphi)$ and define the $\boldsymbol{q}_{{b}}=(\varphi,d)^{\text{T}}$ as virtual joints, where $d=\sqrt{x_{b}^{2}+y_{b}^{2}}$.
For the container fixed at the end for catching the ball, we get the overall position positive kinematic as follows:
${}_{c}^{w}\boldsymbol{T}={}_{b}^{w}\boldsymbol{T}(x_{b},y_{b},\varphi)\cdot{}_{m}^{b}\boldsymbol{T}\cdot{}_{e}^{m}\boldsymbol{T}(\boldsymbol{q}_{{m}})\cdot{}_{c}^{e}\boldsymbol{T}$ (1)
where $w,b,m,e,c$ represent the world, mobile base, manipulator base, end-effector, and container coordinate system, respectively.
Analogous to the Jacobian used in manipulator kinematics, we define the extended Jacobian matrix ${}_{c}^{w}\boldsymbol{J}$ for the mobile manipulator to compute the container’s velocity in the world coordinate system.
${}_{c}^{w}\boldsymbol{v}={}_{c}^{w}\boldsymbol{J}(x_{b},y_{b},\varphi,\boldsymbol{q}_{{m}},{}_{c}^{e}\boldsymbol{T})\dot{\boldsymbol{q}}$ (2)
where $\dot{\boldsymbol{q}}=(\dot{\boldsymbol{q}}_{b},\dot{\boldsymbol{q}}_{m})^{\text{T}}$ and ${}_{c}^{w}\boldsymbol{v}=(v_{x},v_{y},v_{z},\omega_{x},\omega_{y},\omega_{z})^{\text{T}}$. This maps the velocity of whole-body joints to the container’s velocity.
@@
$\ddot{\boldsymbol{p}}=-\boldsymbol{g}-k_{ad}\|\dot{\boldsymbol{p}}\|\dot{\boldsymbol{p}}$ (3)
@@
III-B Problem Formulation
III-B1 Problem 1
Catching a moving ball requires that the container and the ball align in both position and direction. Thus, the goal of catching the ball at the catch time $t_{ca}$ can be achieved by satisfying the following two conditions:
${}_{c}^{w}\boldsymbol{P}(\boldsymbol{q}_{ca})=\boldsymbol{p}(t_{ca})$ (4)
$\frac{{}_{c}^{w}\boldsymbol{z}(\boldsymbol{q}_{ca})}{|{}_{c}^{w}\boldsymbol{z}(\boldsymbol{q}_{ca})|}=-\frac{\dot{\boldsymbol{p}}(t_{ca})}{|\dot{\boldsymbol{p}}(t_{ca})|}$ (5)
where ${}_{c}^{w}\boldsymbol{P}$ denotes the position of the container in the world frame, $\boldsymbol{q}_{ca}$ is the joint configuration of the robot at the catch time $t_{ca}$, ${}_{c}^{w}\boldsymbol{z}(\boldsymbol{q}_{ca})$ indicates the container’s z-axis vector expressed in the world frame.
III-B2 Problem 2
Given the catch configuration $\boldsymbol{q}_{ca}$, we should start with trajectory planning from $q_{0}$. Once the ball is caught, the planning transitions to a compliant control strategy. The two phases, Pre-Catching (PRC) and Post-Catching (POC) can be described as:
$\begin{cases}{\mathit{\Phi}}_{prc}(t;T_{prc},\boldsymbol{q}_{0},\boldsymbol{q}_{ca}),&\quad t<t_{ca}\\ {\mathit{\Phi}}_{poc}(t;\boldsymbol{\boldsymbol{q}},\boldsymbol{\psi}),&\quad t\geq t_{ca}\end{cases}$ (6)
where $T_{prc}$ denotes the trajectory planning time of PRC; $\boldsymbol{\psi}$ represents the output of the compliant control strategy.
III-B3 Problem 3
Given the sequence of ball’s velocity $\boldsymbol{\dot{\boldsymbol{p}}}(t_{ca})$ and position $\boldsymbol{p}(t_{ca})$ before catching, we need to predict the expected compliant sequence $\boldsymbol{\psi}$ of the manipulator to catch the ball with a cushioning mechanism. This can be formulated as follows:
$\boldsymbol{\psi}=\boldsymbol{N}(\boldsymbol{p}(t_{ca}),\boldsymbol{\dot{\boldsymbol{p}}}(t_{ca}))$ (7)
where $\boldsymbol{N}(\cdot{})$ means prediction network.
@@ IV-A Ball Estimation and Prediction
To predict the trajectory of a ball in flight, we first estimate its current state and then use a dynamical model to forecast its future position over a given time horizon. We use a modified discrete-time Kalman filter that accounts for aerodynamic drag, as described in [25]. Given measurements at the discrete time index $k$, the state of the ball at the next time step $\boldsymbol{p}[k+1|k]$ can be predicted as follows:
$\ddot{\boldsymbol{p}}[k|k]=-\boldsymbol{g}-k_{ad}||\dot{\boldsymbol{p}}[k|k]||\dot{\boldsymbol{p}}[k|k]$ (8a)
$\dot{\boldsymbol{p}}[k+1|k]=\dot{\boldsymbol{p}}[k|k]+\delta_{k}\ddot{\boldsymbol{p}}[k|k]$ (8b)
$\boldsymbol{p}[k+1|k]=\boldsymbol{p}[k|k]+{\delta}_{k}\dot{\boldsymbol{p}}[k|k]+\frac{1}{2}{\delta}_{k}^{2}\ddot{\boldsymbol{p}}[k|k],$ (8c)
where $\delta_{k}$ is the interval time between $k$ and $k+1$. The Kalman filter we used follows the standard approach described in [26]. The aerodynamic drag coefficient $k_{ad}$ is estimated offline using a Recursive Least-Squares estimator [25]. Then the ball’s trajectory $\hat{\boldsymbol{p}}$ can be predicted by (3) through numerically integrating over a finite time horizon. We use a cubic spline interpolation between two-time steps to query the $\hat{\boldsymbol{p}}(t)$ and $\hat{\dot{\boldsymbol{p}}}(t)$ at any time.
@@ IV-B High-Level Capture Planner
The high-level planner outputs an optimal catching configuration $\boldsymbol{q}_{ca}$ and catch time $t_{ca}$ that enables the container to intercept the ball’s trajectory at time $t_{ca}$. For this purpose, the following QP problem is formulated:
$\min_{\boldsymbol{q}_{ca},t_{ca}}\frac{1}{2}\left(\boldsymbol{\mathit{\Lambda}}\|\boldsymbol{q}_{ca}-\boldsymbol{q}_{0}\|^{2}-\alpha\|t_{ca}\|^{2}\right)$ (9a)
@@
$\frac{{}_{c}^{w}\boldsymbol{z}(\boldsymbol{q}_{ca})}{|{}_{c}^{w}\boldsymbol{z}(\boldsymbol{q}_{ca})|}=-\frac{\hat{\dot{\boldsymbol{p}}}(t_{ca})}{|\hat{\dot{\boldsymbol{p}}}(t_{ca})|}$ (9c)
${}_{c}^{w}\boldsymbol{T}_{z}(\boldsymbol{q}_{ca})\geq\beta$ (9d)
$\boldsymbol{q}_{ca}^{min}\leq\boldsymbol{q}_{ca}\leq\boldsymbol{q}_{ca}^{max}$ (9e)
@@
The objective function tends to minimize joint movement and extend capture time. $\boldsymbol{\mathit{\Lambda}}$ and $\alpha$ are the weights associated with each aspect. $\boldsymbol{\mathit{\Lambda}}$ can be defined as:
$\boldsymbol{\mathit{\Lambda}}=\begin{pmatrix}\boldsymbol{E}({\mathit{\Lambda}}_{b})&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{E}({\mathit{\Lambda}}_{m})\\ \end{pmatrix}$ (10)
@@
where $\boldsymbol{E}$ represents identity matrix, ${\mathit{\Lambda}}_{b}$, ${\mathit{\Lambda}}_{m}$ denotes the cost values associated with the mobile base and the manipulator. Due to the motion control accuracy of the mobile base is generally lower than the manipulator, ${\mathit{\Lambda}}_{b}$ tends to set higher than ${\mathit{\Lambda}}_{m}$ to encourage more movement of the manipulator.
${}_{c}^{w}\boldsymbol{T}_{z}(\boldsymbol{q}_{ca})$ in (9d) denotes the $Z$-axis height of the container at the capture configuration $\boldsymbol{q}_{ca}$. For the potential collision between the mobile base and the manipulator, [14, 27] use a semi-cylindrical spatial constraint for the position of the container. However, we find that the manipulator typically extends to reach out for a catching. Therefore, by ensuring the height of capture, such a self-collision will also be prevented. Moreover, due to the subsequent PRC phase, it is meaningful for the robot to maintain a specific height when catching the ball. The constraint of ${}_{c}^{w}\boldsymbol{T}_{z}(\boldsymbol{q}_{ca})$ is simple but reasonable and effective.
IV-C Low-Level Joint Planner
IV-C1 PRC Planner

### v3

# CushionCatch: A Compliant Catching Mechanism for Mobile Manipulators via Combined Optimization and Learning

@@ Abstract
Catching flying objects with a cushioning process is a skill commonly performed by humans, yet it remains a significant challenge for robots. In this paper, we present a framework that combines optimization and learning to achieve compliant catching on mobile manipulators (CCMM). First, we propose a high-level capture planner for mobile manipulators (MM) that calculates the optimal capture point and joint configuration. Next, the pre-catching (PRC) planner ensures the robot reaches the target joint configuration as quickly as possible. To learn compliant catching strategies, we propose a network that leverages the strengths of LSTM for capturing temporal dependencies and positional encoding for spatial context (P-LSTM). This network is designed to effectively learn compliant strategies from human demonstrations. Following this, the post-catching (POC) planner tracks the compliant sequence output by the P-LSTM while avoiding potential collisions due to structural differences between humans and robots. We validate the CCMM framework through both simulated and real-world ball-catching scenarios, achieving a success rate of 98.70% in simulation, 92.59% in real-world tests, and a 28.7% reduction in impact torques. The open source code will be released for the reference of the community1.
@@ I INTRODUCTION
Manipulating dynamic objects is challenging, requiring a sophisticated interplay of object detection, motion planning, and control [1]. In dynamic manipulation scenarios, tasks such as throwing [2, 3], juggling [4], and catching [5, 6] are essential. Catching dynamic objects requires quick and precise motion planning, followed by fast and accurate execution. For the motion planning of catching, the process can be divided into two phases, pre-catching (PRC) and post-catching (POC) [7]. In the POC phase, for a high-stiffness robot, the object may bounce away or generate significant impact forces upon contact. To address these problems, incorporating a cushioning mechanism during catching is beneficial, as it reduces impact forces [8, 9], prevents damage to both the robot and the object [10], and minimizes the risk of rebound, especially for high-speed or high-mass objects [11, 12]. Moreover, since humans always exhibit cushioning behavior when catching high-mass objects, an effective approach would be for the robot to learn this motion from humans [7].
Figure 1: The mobile manipulator’s motions under CCMM framework for achieving compliant catching.
Mobile manipulators, with their high degrees of freedom, are better suited to learning from human behavior, enabling them to outperform conventional stationary robots in terms of flexibility and efficiency [13, 14, 15, 16, 17]. Their mobility allows them to actively navigate and interact with objects [18, 19]. In catching tasks, most prior research focused on rigid control approaches, often overlooking the importance of compliant control [1, 6, 14, 20]. We focus on introducing a framework to achieve compliant catching for mobile manipulators (CCMM). Compared with previous works, the contributions of this letter are:
1. A novel CCMM framework effectively addresses all the planning and tracking challenges throughout the entire compliant catching process for mobile manipulators, while being designed for modularity.
2. Propose a P-LSTM network that combines Long Short-Term Memory (LSTM) with positional encoding (PE), designed to effectively learn compliant catching strategies from human demonstrations. Address the potential collisions arising from the structural differences between humans and robots when learning demonstrations.
3. Simulation and physical experiments were conducted on a mobile manipulator to validate the effectiveness of our framework and the role of compliance. A catching success rates of 98.70% in simulation, 92.59% in physical experiments and a 28.7% reduction in impact torque were achieved. An open source code of the proposed method will be released1.
@@ II Related Work
Many studies have explored ball catching tasks using fixed-base manipulators [1, 6, 21, 22], with a particular focus on motion planning for the manipulator. Early research in this area relied on heuristic methods to determine the optimal grasping point and generated motion trajectories through interpolation [22]. However, these approaches often fell short in achieving the most effective capture motion for the robot. A more promising approach involves nonlinear optimization, which helps identify the optimal joint configuration for catching the ball, often in combination with parameterization techniques such as five-order polynomials [1]. The approach in [6] introduces a control law in dynamical systems, allowing the system to perform high-speed manipulations for catching fast-moving objects. However, these methods have typically been limited to fixed-base manipulators.
The work in [20, 23] explores the use of mobile manipulators for catching, considering both the manipulator and the mobile base for coordinated control during capture. However, they overlook the importance of compliance. In reality, a ball thrown from a distance can easily impact a joint, exceeding the torque limit [24]. Current research on cushioning the impact force during capture has focused primarily on fixed-base manipulators [25, 26]. These studies rely on impedance controller based approaches, which are challenging to implement in mobile manipulators for generating unified, coordinated control.
Learning from human demonstrations is an effective method for teaching robots to execute fast and reactive motions. Humans can catch moving objects with nonzero velocity, creating smooth trajectories, even for unexpected objects [27]. In [5], the robot learns motion from human throwing demonstrations, but it halts immediately when the object makes contact with the end-effector. [7] develops nucleated motion primitives from human demonstration data to generate compliant trajectories for catching. However, this research focuses only on the capture of vertically falling objects and not account for the potential safety issues arising from structural differences between humans and robots. Transferring humans’ compliant behavior to mobile manipulators for catching flying objects while ensuring safety, remains an unresolved challenge.
Figure 2: Overview of the system framework: External sensors provide status for both the ball and the robot. The ball’s state is refined using a Kalman filter, which is then used to predict its future trajectory over a time horizon. The high-level capture planner computes the optimal capture joint state, $\bm{q}_{ca}$. The PRC planner generates rapid motion trajectories to $\bm{q}_{ca}$ and sends them to the joint position controller. The PRC planner tracks the compliant velocity sequence from the P-LSTM network, which is designed to learn from human demonstrations.
@@ III-A System Model
The mobile manipulator we use includes a mobile base and a manipulator, with $n=n_{b}+n_{m}$ degrees of freedom. The state of the manipulator is described by joint positions $\bm{q}_{{m}}\in\mathbb{R}^{n_{m}}$. For a mobile base with a differential drive, we use the full location description $(x_{b},y_{b},\varphi)$ and define the $\bm{q}_{{b}}\in\mathbb{R}^{n_{b}}$ as virtual joints.
For the container fixed at the end for catching the ball, we get the overall forward kinematics as follows:
${}_{c}^{w}\bm{T}={}_{b}^{w}\bm{T}(x_{b},y_{b},\varphi)\cdot{}_{m}^{b}\bm{T}\cdot{}_{e}^{m}\bm{T}(\bm{q}_{{m}})\cdot{}_{c}^{e}\bm{T},$ (1)
where $w$ represents the world coordinate system; ${}_{c}^{w}\bm{T}$ represents the homogeneous transformation of the container to the world; ${}_{b}^{w}\bm{T}$ is the transformation of the mobile base frame relative to the world; ${}_{m}^{b}\bm{T}$ denotes a constant relative pose from the mobile base frame to the manipulator frame; ${}_{e}^{m}\bm{T}$ is the position forward kinematics of the manipulator where the end-effector (EE) frame is $e$ and ${}_{c}^{e}\bm{T}$ refers to the transformation from the EE frame to the container frame.
Analogous to the Jacobian used in manipulator kinematics, we define the extended Jacobian matrix ${}_{c}^{w}\bm{J}$ for the mobile manipulator to compute the container’s velocity in the world coordinate system.
${}_{c}^{w}\bm{v}={}_{c}^{w}\bm{J}(x_{b},y_{b},\varphi,\bm{q}_{{m}},{}_{c}^{e}\bm{T})\dot{\bm{q}},$ (2)
where $\dot{\bm{q}}=(\dot{\bm{q}}_{b},\dot{\bm{q}}_{m})^{\text{T}}$ and ${}_{c}^{w}\bm{v}=(v_{x},v_{y},v_{z},\omega_{x},\omega_{y},\omega_{z})^{\text{T}}$. This maps the velocity of whole-body joints to the container’s velocity.
@@
$\ddot{\bm{p}}=-\bm{g}-k_{ad}\|\dot{\bm{p}}\|\dot{\bm{p}},$ (3)
@@ IV-A Ball Estimation and Prediction
To predict the trajectory of a ball in flight, we should first estimate its current state and then use a dynamical model to forecast its future position over a given time horizon. We use a modified discrete-time Kalman filter that accounts for aerodynamic drag, as described in [28]. Given measurements at the discrete time index $k$, the state of the ball at the next time step $\bm{p}[k+1|k]$ can be predicted as follows:
$\ddot{\bm{p}}[k|k]=-\bm{g}-k_{ad}||\dot{\bm{p}}[k|k]||\dot{\bm{p}}[k|k],$ (4a)
$\dot{\bm{p}}[k+1|k]=\dot{\bm{p}}[k|k]+\delta_{k}\ddot{\bm{p}}[k|k],$ (4b)
$\bm{p}[k+1|k]=\bm{p}[k|k]+{\delta}_{k}\dot{\bm{p}}[k|k]+\frac{1}{2}{\delta}_{k}^{2}\ddot{\bm{p}}[k|k],$ (4c)
where $\delta_{k}$ is the time interval between $k$ and $k+1$. The Kalman filter we used follows the standard approach described in [29]. The aerodynamic drag coefficient $k_{ad}$ is estimated offline using a Recursive Least-Squares estimator [28]. Then the ball’s predicted trajectory point $\hat{\bm{p}}$ can be obtained by (3) through numerically integrating over a finite time horizon. We use a cubic spline interpolation between two-time steps to query the $\hat{\bm{p}}(t)$ and $\hat{\dot{\bm{p}}}(t)$ at any time.
@@ IV-B High-Level Capture Planner
The high-level planner outputs an optimal catching configuration $\bm{q}_{ca}\in\mathbb{R}^{n}$ and catch time $t_{ca}\in\mathbb{R}$ that enables the container to intercept the ball’s trajectory at time $t_{ca}$. From current state $\bm{q}_{0}$, catching a moving ball requires the container and ball to be aligned in position and parallel in the direction vector. For this purpose, the following QP problem is formulated:
$\min_{\bm{q}_{ca},t_{ca}}\frac{1}{2}\left(\bm{\mathit{\Lambda}}\|\bm{q}_{ca}-\bm{q}_{0}\|^{2}-\alpha\|t_{ca}\|^{2}\right)$ (5a)
@@
$\frac{{}_{c}^{w}\bm{z}(\bm{q}_{ca})}{|{}_{c}^{w}\bm{z}(\bm{q}_{ca})|}=-\frac{\hat{\dot{\bm{p}}}(t_{ca})}{|\hat{\dot{\bm{p}}}(t_{ca})|},$ (5c)
${}_{c}^{w}\bm{T}_{z}(\bm{q}_{ca})\geq\beta,$ (5d)
$\bm{q}_{ca}^{min}\leq\bm{q}_{ca}\leq\bm{q}_{ca}^{max}.$ (5e)
@@
The objective function tends to minimize joint movement and extend capture time. $\bm{\mathit{\Lambda}}$ and $\alpha$ are the weights associated with each aspect. $\bm{\mathit{\Lambda}}$ can be defined as:
$\bm{\mathit{\Lambda}}=\begin{pmatrix}\bm{E}({\mathit{\Lambda}}_{b})&\bm{0}\\ \bm{0}&\bm{E}({\mathit{\Lambda}}_{m})\\ \end{pmatrix},$ (6)
@@
where $\bm{E}$ represents identity matrix, ${\mathit{\Lambda}}_{b}$, ${\mathit{\Lambda}}_{m}$ denotes the cost values associated with the mobile base and the manipulator. Due to the motion control accuracy of the mobile base is generally lower than the manipulator, ${\mathit{\Lambda}}_{b}$ tends to set higher than ${\mathit{\Lambda}}_{m}$ to encourage more movement of the manipulator.
${}_{c}^{w}\bm{P}\in\mathbb{R}^{3}$ represents the position of the container in the world frame. Equation (5b) ensures that the position of the container is the same as the ball’s position at $\bm{q}_{ca}$ ${}_{c}^{w}\bm{z}(\bm{q}_{ca})\in\mathbb{R}^{3}$ indicates the container’s z-axis vector expressed in the world frame. Through (5c), the direction of the container is opposite to the velocity of the ball at $t_{ca}$, that is, parallel.
${}_{c}^{w}\bm{T}_{z}(\bm{q}_{ca})\in\mathbb{R}$ in (5d) denotes the $Z$-axis height of the container at the capture configuration $\bm{q}_{ca}$. Due to the cushioning motions in subsequent POC phase, (5d) is meaningful for the robot to maintain a specific height when catching the ball.
IV-C PRC Planner